# On Refined Vogel’s universality

### v1

@@ 1 Introduction
    and the symmetric bracket and quantum numbers numbers are defined to be
@@
    The number of variables $x_{i}$ and the length of the Weyl vector coincide and are equal to the dimension of the Euclidean space where the root lattice embeded.
@@ 2 Universality in Chern-Simons theory
Despite the dual quantum dimensions also factorize, it is unclear weather this alternative to quantum dimensions, ${{}^{\vee}\!\hbox{qD}}_{\lambda}^{R}$ can be universalized. It meets, at least, two serious problems. First of all, if there is a universal formula for the dual quantum dimensions, there should exist at once two universal formulas coinciding for the simply laced root systems and differing for the non-simply-laced ones. It is quite an intricate requirement [^3] One can find a discussion of uniqueness of the universal formula for the quantum dimensions in [29].. Second, the dual quantum dimensions for the root systems $B_{n}$ and $C_{n}$ coincide though they are associated with completely distinct Vogel’s parameters with even distinct dependence on the rank $n$. This condition is also not that simple to satisfy.
@@ 3 Universality in refined Chern-Simons theory
$\hbox{Md}^{R}_{\lambda}:=P^{R}_{\lambda}\left(x=q^{2\rho_{k}}\,|\,t_{l}^{2}\,|\,q^{2},t^{2}\right)$ (14)
@@
4 Comments on non-simply-laced algebras
Note that the dual Macdonald dimension in the $B_{n}$ case is equal to that in the $D_{n}$ case at $t_{s}=1$, $\hbox{Md}_{\text{Adj}}^{D_{n}}$ [39]. Moreover, the ordinary Macdonald dimension (16) drastically simplifies and factorizes at $t_{s}=1$ and coincides with $\hbox{Md}_{\text{Adj}}^{D_{n}}$. However, the two root systems $B_{n}$ and $D_{n}$ are described by distinct Vogel’s parameters, hence, the $B_{n}$ series still does not match the universality formula at $t_{s}=1$. At last, note the Macdonald dimension in the $B_{n}$ case (16) at $t_{s}=q$ also factorizes and takes an intriguing form
@@ 4 Comments on non-simply-laced algebras
The technical reason for the $C_{n}$ case to look more complicated is that, in the minimal normalization (which is associated with the Vogel’s universality), the answers are far more complicated [^5] For instance, a counterpart of formula (29) in the standard normalization, when the long root of the root system $C_{n}$ has length 2, is $\overline{P}^{C_{n}}_{\text{\text{Adj}}}\left(x=q^{2\rho_{k}}\,|\,t_{l}^{2}\,|\,q^{2},t^{2}\right)\Big{|}_{t_{l}=q}={\{t^{n}\}\{qt^{n}\}\{q^{3}t^{2n-2}\}\{q^{6}t^{2n-2}\}\over\{q^{2}t^{n-1}\}\{q^{3}t^{n-1}\}\{t\}\{qt\}}$ (30) and one can see, it is again factorized at the point $t_{l}=q$ as in the $B_{n}$ case., and the minimal normalization makes a difference only in the $C_{n}$ and $G_{2}$ cases.

### v2

@@ 1 Introduction
    and the symmetric bracket and quantum numbers are defined to be
@@
    The number of variables $x_{i}$ and the length of the Weyl vector coincide and are equal to the dimension of the Euclidean space where the root lattice embedded.
@@ 2 Universality in Chern-Simons theory
Despite the dual quantum dimensions also factorize, it is unclear weather this alternative to quantum dimensions, ${{}^{\vee}\!\hbox{qD}}_{\lambda}^{R}$ can be universalized. It meets, at least, two serious problems. First of all, if there is a universal formula for the dual quantum dimensions, there should exist at once two universal formulas coinciding for the simply laced root systems and differing for the non-simply-laced ones. It is quite an intricate requirement [^3] One can find a discussion of uniqueness of the universal formula for the quantum dimensions in [29].. Second, the dual quantum dimensions for the root systems $B_{n}$ and $C_{n}$ coincide (which is a corollary of duality between the symplectic and orthogonal groups [30, 31, 32, 33]) though they are associated with completely distinct Vogel’s parameters with even distinct dependence on the rank $n$. This condition is also not that simple to satisfy.
@@ 3 Universality in refined Chern-Simons theory
$\hbox{Md}^{R}_{\lambda}:=P^{R}_{\lambda}\left(x=q^{2\rho_{k}}\,|\,t_{\alpha}^{2}\,|\,q^{2},t^{2}\right)$ (14)
@@
4 Comments on non-simply-laced root systems
Note that the dual Macdonald dimension in the $B_{n}$ case is equal to that in the $D_{n}$ case at $t_{s}=1$, $\hbox{Md}_{\text{Adj}}^{D_{n}}$ [43]. Moreover, the ordinary Macdonald dimension (16) drastically simplifies and factorizes at $t_{s}=1$ and coincides with $\hbox{Md}_{\text{Adj}}^{D_{n}}$. However, the two root systems $B_{n}$ and $D_{n}$ are described by distinct Vogel’s parameters, hence, the $B_{n}$ series still does not match the universality formula at $t_{s}=1$.
The reason for coinciding dimensions is that there are duality relations [43, 42] (the polynomials here are evaluated at the normalization [^5] We remind that the dual Macdonald dimensions do not depend on normalization. where for all roots but may be one $(\alpha,\alpha)$=2)
[…]
In fact, these relations can be extended to other representations (note that $[1,1]$ is the adjoint representation for $B_{n}$ and $D_{n}$, and $[2]$, the adjoint representation for $C_{n}$). However, for constructing universality formulas, this kind of relations is of no use since the Macdonald polynomials at $t_{s,l}=1$ do not match at $t=q$ the corresponding Lie group characters and, hence, the quantum dimensions.
At last, note the Macdonald dimension in the $B_{n}$ case (16) at $t_{s}=q$ also factorizes and takes an intriguing form
@@ 4 Comments on non-simply-laced root systems
The technical reason for the $C_{n}$ case to look more complicated is that, in the minimal normalization (which is associated with the Vogel’s universality), the answers are far more complicated [^6] For instance, a counterpart of formula (30) in the standard normalization, when, for the long root of the root system $C_{n}$, $(\alpha_{l},\alpha_{l})=4$ is $\overline{P}^{C_{n}}_{\text{\text{Adj}}}\left(x=q^{2\rho_{k}}\,|\,t_{l}^{2}\,|\,q^{2},t^{2}\right)\Big{|}_{t_{l}=q}={\{t^{n}\}\{qt^{n}\}\{q^{3}t^{2n-2}\}\{q^{6}t^{2n-2}\}\over\{q^{2}t^{n-1}\}\{q^{3}t^{n-1}\}\{t\}\{qt\}}$ (31) and one can see, it is again factorized at the point $t_{l}=q$ as in the $B_{n}$ case., and the minimal normalization makes a difference only in the $C_{n}$ and $G_{2}$ cases.